What is the average velocity of a man who ran a distance of 3km at a

velocity of 7.5 m/s, and the walked 1km at a velocity of 0.8 m/s?



1
Expert's answer
2021-02-07T19:19:57-0500

Average velocity can be found as follows:


"v_{avg}=\dfrac{Total\ displacement}{Total\ time\ taken}=\dfrac{d_{tot}}{t_{tot}}."

Let's find the total displacement:


"d_{tot}=d_1+d_2=3000\ m+1000\ m=4000\ m."

Let's find the total time taken:


"t_{tot}=\dfrac{d_1}{v_1}+\dfrac{d_2}{v_2},""t_{tot}=\dfrac{3000\ m}{7.5\ \dfrac{m}{s}}+\dfrac{1000\ m}{0.8\ \dfrac{m}{s}}=1650\ s."

Finally, we can find the average velocity:


"v_{avg}=\dfrac{4000\ m}{1650\ s}=2.42\ \dfrac{m}{s}."